# PSHuman: Photorealistic Single-view Human Reconstruction using Cross-Scale Diffusion

### v1

@@ Abstract
Detailed and photorealistic 3D human modeling is essential for various applications and has seen tremendous progress. However, full-body reconstruction from a monocular RGB image remains challenging due to the ill-posed nature of the problem and sophisticated clothing topology with self-occlusions. In this paper, we propose PSHuman, a novel framework that explicitly reconstructs human meshes utilizing priors from the multi-view diffusion model. It is found that directly applying multiview diffusion on single-view human images leads to severe geometric distortions, especially on generated faces. To address it, we propose a cross-scale diffusion that models the joint probability distribution of global full-body shape and local facial characteristics, enabling detailed and identity-preserved novel-view generation without any geometric distortion. Moreover, to enhance cross-view body shape consistency of varied human poses, we condition the generative model on parametric models like SMPL-X, which provide body priors and prevent unnatural views inconsistent with human anatomy. Leveraging the generated multi-view normal and color images, we present SMPLX-initialized explicit human carving to recover realistic textured human meshes efficiently. Extensive experimental results and quantitative evaluations on CAPE and THuman2.1 datasets demonstrate PSHuman’s superiority in geometry details, texture fidelity, and generalization capability.
Figure 1: We introduce PSHuman, a diffusion-based full-body human reconstruction model. Given a single image of a clothed person, our method facilitates detailed geometry and realistic 3D human appearance across various poses within one minute.
@@ 1 Introduction
Photorealistic 3D reconstruction of clothed humans is a promising and widely investigated research domain with significant applications across several industries, including gaming, movies, fashion, and virtual/augmented reality (AR/VR) (Ma et al., 2021; Orts-Escolano et al., 2016). Traditional methods, which perform multiview stereo and non-rigid registration using multi-camera setups or incorporate additional depth signals, have achieved accurate modeling. However, reconstruction from an in-the-wild RGB image remains an open problem due to sophisticated body poses and complex clothing topology.
A plethora of studies have been developed to address these challenges. PIFu (Saito et al., 2019) and related efforts (Saito et al., 2020; Zhang et al., 2024b; Ho et al., 2024; Zhang et al., 2024a; Xiu et al., 2022) extract pixel-aligned features from the color or normal image and leverage implicit functions to predict the occupancy field (Mescheder et al., 2019) of the 3D human body and ECON (Xiu et al., 2023) utilizes bilateral normal integration (BiNI) to lift normal clues to 3D body to remain predicted details explicitly. On the one hand, these efforts indeed lead to improvements in terms of either monocular ambiguity or postural intricacy through the introduction of other geometric clues or occluded-view information. On the other hand, the direct regression paradigm still falls short in detail loss and artifacts. Similarly, recent progress in appearance reconstruction (Zhang et al., 2024b; Ho et al., 2024) follows the implicit function to infer full-body texture, struggling with texture unrealism due to poor generalization capability.
  Input    PaMIR     PIFuHD       GTA        SiFU       ECON         Ours
Figure 2: Geometry comparison between Implicit and Explicit methods.
In this study, we aim to tackle these existing challenges by introducing a multiview diffusion model and a normal-guided explicit human reconstruction framework. We build upon the recent progress of diffusion-based multiview generation models to explore their hallucination capabilities for robust human modeling. As depicted in Fig. 4, PSHuman takes a full-body human image as input, followed by a carefully designed multiview diffusion model and an SMPLX-initialized mesh carving module, outputting a textured 3D human mesh.
@@
Specifically, we fine-tune a pre-trained text-to-image diffusion model (such as Stable Diffusion (Rombach et al., 2022b)) to generate multiview color and normal maps conditioned on the input reference. Despite impressive generative performance, this base framework faces two major challenges: 1) Unnatural body structures, where diffusion models struggle to generate reasonable novel views of posed humans, often resulting in disproportionate body proportions or missing body parts. This issue arises from the severe self-occlusion in the posed human image and lack of body prior for generative models. To address this, we propose an SMPL-X conditioned diffusion model, which concatenates renderings of estimated SMPL-X with the input image to provide pose guidance for novel-view generation. This approach constrains the diffusion model to generate consistent views that adhere to human anatomy, even when fine-tuning with as few as $3,000$ human scans. 2) Face distortion, where pre-trained diffusion models often produce distorted and unnatural face details, especially for full-body human input. This problem is attributed to the small size of the face in full-body images, which provides limited information for detailed normal prediction after VAE encoding. To accurately recover face geometry, we propose a body-face cross-scale diffusion framework that simultaneously generates multiview full-body images and local face ones. We also employ a simple yet efficient noise blending layer to enhance face details in global image, guaranteeing both cross-scale and cross-view consistency. Consequently, PSHuman generates high-quality and detailed novel-view human images and corresponding normal maps.
@@
To fully leverage the generated multiview images, we present an SMPLX-initialized explicit human carving module for fast and high-fidelity textured human mesh modeling. Unlike implicit functions that use Multilayer Perceptrons (MLPs) to map normal features to an implicit surface, or BiNI (Cao et al., 2022) that utilizes variational normal integration to recover 2.5D surfaces, we directly reconstruct the 3D mesh supervised by generated multiview normal maps. In practice, we initialize the human model with predicted SMPL-X, and deform and remesh it with differentiable rasterization Palfinger (2022). As shown in Fig. 2, PSHuman can preserve fine-grained details, such as facial features and fabric wrinkles, and generate natural and harmonious novel views. For texturing on the generated meshes, we first fuse multiview color images using differentiable rendering to mitigate generative inconsistencies, then project them onto the reconstructed 3D mesh.
@@
The entire reconstruction process takes as few as one minute. It is noted that recent SDS-based methods (Huang et al., 2024b; a) also achieve state-of-the-art performance in geometry details and appearance fidelity. However, they can only handle simple poses and suffer from time-consuming optimization (such as TeCH Huang et al. (2024b), which takes approximately six hours). Conversely, PSHuman achieves a balance between precision, efficiency, and pose robustness.
@@
  • We present a body-face cross-scale diffusion and an SMPL-X conditioned multi-view diffusion for high-quality full-body human image generation with high-fidelity face details.
@@
  • We design an SMPLX-initialized explicit human carving module to fast recover textured human mesh based on generated multi-view cross-domain images, achieving SOTA performance on THuman2.1 and CAPE datasets.
2 Related Work
Single-image human reconstruction has seen rapid advancements in recent years, primarily driven by three key approaches: implicit function-based reconstruction, explicit shape-based reconstruction, and the emerging 2D diffusion-based methods.
@@ 2 Related Work
Implicit Human Reconstruction. Implicit functions have gained significant traction in human reconstruction (Chibane et al., 2020; Gropp et al., 2020; Yang et al., 2023) due to their flexibility in handling complex topology and diverse clothing styles. Pioneering works such as PIFu Saito et al. (2019) introduce pixel-aligned implicit functions, mapping 2D image features to 3D implicit surface for continuous modeling. Building upon this, subsequent research incorporates parametric models (e.g., SMPL) to enhance anatomical plausibility and robustness in challenging in-the-wild poses (He et al., 2020; Xiu et al., 2022; Zheng et al., 2021; Zhang et al., 2024a) or for animation-ready modeling (Huang et al., 2020; He et al., 2021). Other efforts enhance geometric details and dynamic stability by introducing normal (Saito et al., 2020), depth clues  (Yu et al., 2021b; Zheng et al., 2023), or decoupling albedo (Alldieck et al., 2022) from natural inputs. However, these methods struggle with unseen areas due to limited observed information. More recent approaches (Zhang et al., 2024b; Ho et al., 2024) incorporate predicted side-view images to enhance visualization but still face challenges in balancing quality, efficiency, and robustness.
@@
Explicit Human Reconstruction. Early research focuses on explicit representation for human reconstruction. Voxel-based methods (Varol et al., 2018; Zheng et al., 2019) utilize 3D UNet to predict volumetric confidence occupied by the human body, which demands high memory and often results in compromised spatial resolution, hindering the capture of fine details crucial for realistic representation. As a more efficient alternative, visual hulls (Natsume et al., 2019) approximate 3D shapes by incorporating silhouettes and 3D joints. Another strategy involves using depth  (Gabeur et al., 2019; Smith et al., 2019; Han et al., 2023) or normal (Alldieck et al., 2019; Xiu et al., 2023) information to explicitly infer the 3D human body, balancing detail preservation with computational efficiency. Among these, ECON utilizes normal integration and shape completion, achieving extreme robustness for challenging poses and loose clothing. The major limitations lie in sub-optimal geometry and supporting appearance. To address this, we propose to simultaneously recover geometry and appearance with differentiable rasterization under the supervision of multi-view normal and color maps predicted by the diffusion model.
@@
Diffusion-based Human Reconstruction. Most recently, Score Distillation Sampling (SDS) Poole et al. (2022) based human generation methods (Liao et al., 2023; Huang et al., 2024b) have achieved SOTA performance. However, these approaches often require time-consuming optimization. Draw inspiration from the advancement of multi-view diffusion based 3D generation (Liu et al., 2023; Long et al., 2024; Li et al., 2024; Voleti et al., 2024; Tang et al., 2024), our work reduces the inference time by directly generating multiple human views for human reconstruction. We further augment human generation capabilities through the introduction of a novel SMPL-X-conditioned cross-scale attention framework. Most related to our work, Chupa Kim et al. (2023) also reconstructs with multi-view normals. However, it still depends on optimization-based refinement and does not support image condition and texture modeling.
3 Our Approach
@@
Overview. Given a single color image, we aim to reconstruct the textured 3D human mesh with generated realistic invisible views. PSHuman is built upon recent multi-view generative models (Li et al., 2024; Long et al., 2024), including two primary stages: 1) a body-face cross-scale diffusion model conditioned on SMPL-X, which generates multi-view full-body cross-domain (color and normal) images and local facial ones (Sec. 3.1), 2) an SMPLX-initialized explicit human carving module for modeling 3D textured meshes (Sec. 3.2). Since we generate normal maps and images, we use $x$ and $z$ as the raw data and latents for both data modalities.
@@
Figure 4: Overall pipeline. Given a single full-body human image, PSHuman recovers the texture human mesh by two stages: 1) Body-face enhanced and SMPL-X conditioned multi-view generation. The input image and predicted SMPL-X are fed into a multi-view image diffusion model to generate six views of global full-body images and local face images. 2) SMPLX-initialized explicit human carving. Utilizing generated normal and color maps to deform and remesh the SMPL-X with differentiable rasterization.
3.1 Body-face multi-view diffusion
@@ 3.1.1 Body-Face diffusion
Motivation. Simply adopting the multiview diffusion (Li et al., 2024; Long et al., 2024) for 3D human reconstruction leads to distorted faces and changes of face identities in the reconstruction results. Because the face only occupies a small region with a low resolution in the image and cannot be accurately generated by the multiview diffusion model. Since humans are very sensitive to slight changes in faces, such generation inaccuracy of faces leads to obvious distortion and identity changes. This motivates us to separately apply another multiview diffusion model to generate the face at a high resolution with more accuracy.
@@
where $q$ defines the forward process to add noises to the original data and $p$ defines the reverse process to generate data by denoising. For the forward process, we simply omit the condition on the $x^{F}_{t-1}$ and add noises to the face and body images separately by the approximated forward process
@@
Although explicitly defining forward process for $q(x^{B}_{t}|x^{B}_{t-1},x^{F}_{t-1})$ is feasible for the vanilla diffusion model, it is difficult for the latent diffusion model. We explain this difficulty and the feasibility of this approximation in Sec. A.1. For the reverse process $p(x_{t-1}|x_{t})$, the face diffusion is just a vanilla diffusion model $p(x^{F}_{t-1}|p^{F}_{t})$ while the body diffusion model will additionally use the face denoising results as conditions by $p(x^{B}_{t-1}|p^{B}_{t},p^{F}_{t-1})$, as shown in Fig. 5, which is implemented by the following joint denoising scheme.
Figure 5: Illustration of joint denoising diffusion block.
Joint denoising. We utilize a simple but efficient noise blending layer to jointly denoise in body-face diffusion. Specifically, in each self-attention block of UNet, we extract the latent vector of the face branch, resize it with scale $s$, and add it to the face region of the global branch with a weighted sum. Specifically, let us take one of the hidden layers as an example. We denote $h^{B_{n}}_{t}$ and $h^{F}_{t}$ as hidden vectors of the $n$-th body view and face view at the same attention layer  [^1] Here, we omit the layer subscript for simplicity. and timestep $t$, the blending operation can be written as
@@
where, the $\mathop{RP}$ is the resize and padding function, $w$ is the mask of the face region. The resulting latent vector can be represented by $z^{B_{n}}_{t}$ and $z^{F}_{t}$. We jointly optimize the body and face distribution with the following loss,
$\ell=\mathbb{E}_{t,\mathbf{z}^{F}_{0},\mathbf{\epsilon}}\left[\|\mathbf{\epsilon}-\mathbf{\epsilon}_{\theta}(z^{F}_{t},t)\|_{2}\right]+\mathbb{E}_{t,\mathbf{z}^{B}_{0},\mathbf{z}^{F}_{0},n,\mathbf{\epsilon}}\left[\|\mathbf{\epsilon}^{(n)}-\mathbf{\epsilon}^{(n)}_{\theta}(\mathbf{z}^{B}_{t},\mathbf{z}^{F}_{t},t)\|_{2}\right],$ (6)
where $\theta$ is shared weights between face and multiple body views. The noise blending allows the face information to be transferred to novel body views with cross-view attention, improving the overall consistency of generated human images.
3.1.2 SMPL-X guided multi-view diffusion
The diffusion model excels in generating plausible novel views for simple, non-occluded body poses, producing natural human geometry. However, it faces significant challenges with in-the-wild images that often feature self-occlusions. These occlusions can lead to “hallucinations” that violate human structural integrity or exhibit inconsistent limb poses. For example, Fig. 3 illustrates two common issues: (a) the model generating upright side views for a bending posture input, and (b) inconsistencies in arm regions of side views due to self-occlusion, resulting in failed reconstruction.
To mitigate these impediments, we propose incorporating additional pose guidance into the diffusion process. Our method first estimates the SMPL-X parameters of the input image and renders them from six target viewpoints. We then utilize a pre-trained Variational Autoencoder (VAE) encoder to convert these renderings into latent vectors, which are concatenated with noise samples and the reference image to serve as input of the denoising UNet. The introduction of these conditional signals constrains the multi-view distribution, leading to more accurate and consistent human image generation. This approach significantly enhances the model’s generalization capability on complex human poses with self-occlusion.

### v2

# PSHuman: Photorealistic Single-image 3D Human Reconstruction using Cross-Scale Multiview Diffusion and Explicit Remeshing

@@ Abstract
Photorealistic 3D human modeling is essential for various applications and has seen tremendous progress. However, existing methods for monocular full-body reconstruction, typically relying on front and/or predicted back view, still struggle with satisfactory performance due to the ill-posed nature of the problem and sophisticated self-occlusions. In this paper, we propose PSHuman, a novel framework that explicitly reconstructs human meshes utilizing priors from the multiview diffusion model. It is found that directly applying multiview diffusion on single-view human images leads to severe geometric distortions, especially on generated faces. To address it, we propose a cross-scale diffusion that models the joint probability distribution of global full-body shape and local facial characteristics, enabling identity-preserved novel-view generation without geometric distortion. Moreover, to enhance cross-view body shape consistency of varied human poses, we condition the generative model on parametric models (SMPL-X), which provide body priors and prevent unnatural views inconsistent with human anatomy. Leveraging the generated multiview normal and color images, we present SMPLX-initialized explicit human carving to recover realistic textured human meshes efficiently. Extensive experiments on CAPE and THuman2.1 demonstrate PSHuman’s superiority in geometry details, texture fidelity, and generalization capability.
@@ 1 Introduction
Photorealistic 3D reconstruction of clothed humans is a promising and widely investigated research domain with significant applications across several industries, including gaming, movies, fashion, and AR/VR [26, 29]. Traditional methods, which perform multiview stereo and non-rigid registration using multi-camera setups or incorporate additional depth signals, have achieved accurate modeling. However, reconstruction from an in-the-wild RGB image remains an open problem due to sophisticated body poses and complex clothing topology.
@@
Early studies [35, 36, 42] utilize implicit functions [27, 31] to recover textured human mesh from a single color or normal image. Despite improvements in monocular ambiguity and postural intricacy, this regression paradigm still falls short in detail loss and novel view artifacts. Recent efforts [51, 13] incorporate generative information, such as predicting a back view, to mitigate these issues. On the one hand, the reconstruction pipelines still follow implicit functions, which exhibit limitations in capturing high-fidelity geometry and texture details. On the other hand, the introduction of the back view fails to provide enough stereo information to mitigate spatial ambiguity.
In this study, we aim to tackle these existing challenges by introducing a multiview diffusion model and a normal-guided explicit human reconstruction framework. Different from the front and/or back views by existing methods [43, 13], we explore the direct multiple views generation for robust human modeling. As depicted in Fig. 2, PSHuman takes a full-body human image as input, followed by a designed multiview diffusion model and an SMPLX-initialized mesh carving module, outputting a textured 3D human mesh.
@@
Specifically, we fine-tune a pre-trained text-to-image diffusion model (such as Stable Diffusion [34]) to generate multiview color and normal maps conditioned on the input reference. Despite impressive generative performance, this base framework faces two major challenges: 1) Unnatural body structures, where diffusion models struggle to generate reasonable novel views of posed human, often resulting in disproportionate body proportions or missing body parts. As shown in Fig. 1, this issue arises from the severe self-occlusion in the posed human image and lack of body prior for generative models. To address this, we propose a SMPL-X conditioned diffusion model, which concatenates renderings of estimated SMPL-X with the input image to provide pose guidance for novel-view generation. This approach constrains the diffusion model to generate consistent views that adhere to human anatomy, even when fine-tuning with as few as $3,000$ human scans. 2) Face distortion, where pre-trained diffusion models often produce distorted and unnatural face details, especially for full-body human input. This problem is attributed to the small size of the face in full-body images, which provides limited information for detailed normal prediction after VAE encoding. To accurately recover face geometry, we propose a body-face cross-scale diffusion framework that simultaneously generates multiview full-body images and local face ones. We also employ a simple yet efficient noise blending layer to enhance face details in global image, guaranteeing both cross-scale and cross-view consistency. Consequently, PSHuman generates high-quality and detailed novel-view human images and corresponding normal maps.
@@
To fully leverage the generated multiview images, we present a SMPLX-initialized explicit human carving module for fast and high-fidelity textured human mesh modeling. Unlike implicit functions that use Multilayer Perceptrons (MLPs) to map normal features to an implicit surface, or BiNI [3] that utilizes variational normal integration to recover 2.5D surfaces, we directly reconstruct the 3D mesh supervised by generated multiview normal maps. In practice, we initialize the human model with predicted SMPL-X, and deform and remesh it with differentiable rasterization [30]. As shown in Fig. LABEL:fig_teaser, PSHuman can preserve fine-grained details, such as facial features and fabric wrinkles, and generate natural and harmonious novel views. For texturing on the generated meshes, we first fuse multiview color images using differentiable rendering to mitigate generative inconsistencies, then project them onto the reconstructed 3D mesh.
@@
The entire reconstruction process takes as short as one minute. It is noted that recent SDS-based methods [15, 14] also achieve state-of-the-art performance in geometry details and appearance fidelity. However, they can only handle simple poses and suffer from time-consuming optimization (e.g., TeCH [15] takes approximately six hours). Conversely, PSHuman achieves a balance between precision, efficiency, and pose robustness.
@@
  • We present a body-face cross-scale diffusion and a SMPL-X conditioned multiview diffusion for high-quality full-body human image generation with high-fidelity face details.
@@
  • We design a SMPLX-initialized explicit human carving module to fast recover textured human mesh based on generated multiview cross-domain images, achieving SOTA performance on THuman2.1 and CAPE datasets.
2 Related Works
@@
Implicit Human Reconstruction. Implicit functions have gained significant traction in human reconstruction [4, 8, 44] due to their flexibility in handling complex topology and diverse clothing styles. Pioneering works such as PIFu [35] introduce pixel-aligned implicit functions, mapping 2D image features to 3D implicit surface for continuous modeling. Building upon this, subsequent research incorporates parametric models (e.g., SMPL) to enhance anatomical plausibility and robustness in challenging in-the-wild poses [10, 42, 54, 50] or for animation-ready modeling [16, 11]. Other efforts enhance geometric details and dynamic stability by introducing normal [36], depth clues  [47, 52], or decoupling albedo [2] from natural inputs. However, these methods struggle with unseen areas due to limited observed information. More recent approaches [51, 13] incorporate predicted side-view images to enhance visualization but still face challenges in balancing quality, efficiency, and robustness.
@@
Explicit Human Reconstruction. Early research focuses on explicit representation for human reconstruction. Voxel-based methods [39, 53] utilize 3D UNet to predict volumetric confidence occupied by the human body, which demands high memory and often results in compromised spatial resolution, hindering the capture of fine details crucial for realistic representation. As a more efficient alternative, visual hulls [28] approximate 3D shapes by incorporating silhouettes and 3D joints. Another strategy involves using depth  [6, 37, 9] or normal [1, 43] information to explicitly infer the 3D human body, balancing detail preservation with computational efficiency. Among these, ECON utilizes normal integration and shape completion, achieving extreme robustness for challenging poses and loose clothing. The major limitations lie in sub-optimal geometry and supporting appearance. To address this, we propose to simultaneously recover geometry and appearance with differentiable rasterization under the supervision of multiview normal and color maps predicted by the diffusion model.
@@
Diffusion-based Human Reconstruction. Most recently, Score Distillation Sampling (SDS) [32] based human generation methods [22, 15] have achieved SOTA performance. However, these approaches often require time-consuming optimization. Drawing inspiration from the advancement of multiview diffusion based 3D generation [23, 24, 21, 40, 38], our work reduces the inference time by directly generating multiple human views for human reconstruction. We further augment human generation capabilities through the introduction of a novel SMPL-X-conditioned cross-scale attention framework. Most related to our work, Chupa [19] also reconstructs with multiview normals. However, it still depends on optimization-based refinement and does not support image condition and texture modeling.
3 Method
@@
Overview. Given a single color image, PSHuman recovers a textured human mesh by two primary stages: 1) a body-face cross-scale multiview diffusion conditioned on SMPL-X, which generates multiview full-body cross-domain (color and normal) images and local facial ones (Sec. 3.1), 2) an SMPLX-initialized explicit human carving module for modeling 3D textured meshes (Sec. 3.2). Different from previous works utilizing front and/or back views, we follow [21, 24] to directly generate six views (front, front left, left, back, right, and front right) for explicit reconstruction, which strike the best balance between computational cost and effectiveness. Since we generate normal maps and images, we use $x$ and $z$ as the raw data and latent for both modalities.
@@
Figure 2: (a) Overall pipeline. Given a single full-body human image, PSHuman recovers the texture human mesh by two stages: 1) Body-face enhanced and SMPL-X conditioned multiview generation. The input image and predicted SMPL-X are fed into a multiview image diffusion model to generate six views of global full-body images and front local face images. 2) SMPLX-initialized explicit human carving. Utilizing generated normal and color maps to deform and remesh the SMPL-X with differentiable rasterization. (b) Illustration of joint denoising diffusion block.
3.1 Body-face Multiview Diffusion
@@ 3.1.1 Body-face Diffusion
Motivation. Simply adopting the multiview diffusion [21, 24] for 3D human reconstruction leads to distorted faces and altered facial identities. Because the face only occupies a small region with a low resolution in the image and cannot be accurately generated by the multiview diffusion model. Since humans are very sensitive to slight changes in faces, such generation inaccuracy of faces leads to obvious distortion and identity changes. This motivates us to separately apply another multiview diffusion model to generate the face at a high resolution with more accuracy.
@@
where $q$ defines the forward process to add noise to the original data and $p$ defines the reverse process to generate data by denoising. For the forward process, we omit the condition on the $x^{F}_{t-1}$ and add noises to the face and body images separately by the approximated forward process
@@
Although explicitly defining forward process for $q(x^{B}_{t}|x^{B}_{t-1},x^{F}_{t-1})$ is feasible for the vanilla diffusion model, it is difficult for the latent diffusion model. We explain this difficulty and the feasibility of this approximation in supplementary material. For the reverse process $p(x_{t-1}|x_{t})$, the face diffusion is just a vanilla diffusion model $p(x^{F}_{t-1}|p^{F}_{t})$ while the body diffusion model will additionally use the face denoising results as conditions by $p(x^{B}_{t-1}|p^{B}_{t},p^{F}_{t-1})$, as shown in Fig. 2(b), which is implemented by the following joint denoising scheme.
Joint denoising. We utilize a simple but efficient noise blending layer to jointly denoise in body-face diffusion. Specifically, in each self-attention block of UNet, we extract the latent vector of the face branch, resize it with scale $s$, and add it to the face region of the global branch with a weight $w$. Specifically, let us take one of the hidden layers as an example. We denote $h^{B_{n}}_{t}$ and $h^{F}_{t}$ as hidden vectors of the $n$-th body view and face view at the same attention layer  [^1] Here, we omit the layer subscript for simplicity. and timestep $t$, the blending operation can be written as
@@
where the $\mathop{RP}$ is the resize and padding function, and $w$ is a binary mask of the face region, which is obtained with a face detector or a straightforward cropping strategy. The resulting latent vector can be represented by $z^{B_{n}}_{t}$ and $z^{F}_{t}$. We jointly optimize the body and face distribution with the following loss,
$\ell=$ $~{}~{}~{}~{}\mathbb{E}_{t,{z}^{F}_{0},\mathbf{\epsilon}}\left[\|\mathbf{\epsilon}-\mathbf{\epsilon}_{\theta}(z^{F}_{t},t)\|_{2}\right]$
$\ \ +\mathbb{E}_{t,{z}^{B}_{0},{z}^{F}_{0},n,\mathbf{\epsilon}}\left[\|\mathbf{\epsilon}^{(n)}-\mathbf{\epsilon}^{(n)}_{\theta}(z^{B}_{t},z^{F}_{t},t)\|_{2}\right],$ (6)
where $\theta$ is shared weights between face and body views. The noise blending allows the face information to be transferred to novel body views with cross-view attention, improving the overall consistency of generated human images.
3.1.2 SMPL-X Guided Multiview Diffusion
Our multiview diffusion model excels in generating plausible novel views for simple posed images, producing natural human geometry. However, it faces challenges with in-the-wild images that often feature self-occlusions. These occlusions can lead to “hallucinations” that violate human structural integrity or exhibit inconsistent limb poses. For example, Fig. 1 illustrates two common issues: (a) the model generating upright side views for a bending posture, and (b) inconsistencies in arm regions of side views due to self-occlusion, resulting in failed reconstruction.
To mitigate these impediments, we propose incorporating additional pose guidance into the diffusion process. Our method first estimates the SMPL-X of the input image and renders them from six target viewpoints. We then utilize a pre-trained Variational Autoencoder (VAE) encoder to convert SMPL-X renderings and reference images into latent vectors, which are concatenated with noise samples to serve as input of the denoising UNet. The introduction of these conditional signals constrains the multiview distribution, leading to more accurate and consistent human image generation. This approach significantly enhances the model’s generalization capability on complex human poses with self-occlusion.